Find thevenin Equiv. Q23 by R. R1 R2 32 nu
Answer #1

33 Thevenin voltage (Beated CVTA) 38 A 3A VER 1. currents of Let is, is be the mesh meshes 0, © respectively. - current i = 3CVL Apply Kve for mest ♡ : -5 +1 (91) +2 (,-2) =0 substitute iz = 3A =) -5+ i; + 21, -2(3) =0 . 3°, - 576 - , -www 32 - Apply KVL as shown wn ab above for VT: O 2(12-;;) +3 iz + VTT so =) 2(3-ų) +3(3) + VT6 = 0> Vith = -2 (-3) -9 Vin = 4-27 = -23 . Vih= -7.667 volts Thevenin Resistance Rith - Rith is the equivalent Resistance acrossin and 21 are in paollel . Resultant = 1 (2) = 2/22 3 2 1+2 nu 32 2 / 3 ERR .: R15 = +3 = 21 i Rith = 3.667 2 The venin Equiv